If the equation for market demand curve is qd=100-0.5p and the equation for a market supply curve is Qs=-20+p, the market equilibrium price is?


1
Expert's answer
2021-06-03T18:35:02-0400

Solution:

At equilibrium: Quantity demanded (Qd) = Quantity supplied (Qs)

Quantity demanded = 100 – 5p

Quantity supplied = -20 + p

Therefore:

100 – 5p = -20 + p

100 + 20 = p + 5p

120 = 6p


P = "\frac{120}{6} = 20"

The market equilibrium price (p) = 20